4 September 2009 LBC Checkout

Observer: O. Kuhn
Telescope Operator: D. Gonzalez Huerta

Test L2 position

Check Focus (z-SLOAN)

Install and test new mcspu

Running with TCS Build 29 now (same as June/July)

22:30 Opened after some clouds earlier

22:55 LBC was on but Red channel not reporting Vacuum or Temperature. Turned off/on LBC & HK. In doing so, I got an error with the Red rotator, and Red rotator & filter ended in "parking" (not "enabled") state. Ran lbckill/lbcstart to clear UI and tried turning on again. First try, only blue HK came up. Second try - OK. All up and running now.

23:00 taking 10 biases.

23:10 running dofpia (NOTE: need to ssh -Y and ssh -X to display from station in polycom remote observing room --- I don't need to do this from my desktop. An ssh -Y alias for LBTO account on du16 would be helpful).
Note from JMH the following day: I also don't have to do this from my desktop, but I am unable to find anything different between my ssh setup and the LBTO ssh setup. I'm starting to think that something in IDL is set differently, since TCS GUIs display fine for LBTO.

23:17 Red pupils were very tiny and dofpia exited with an error:
POLY_FIT: Warning: Invert detected a small pivot element. 
Rejected pupil with diameter 8.92384
I'm not sure how it got to this state - cleared active optics at start of night. DX and SX collimation tables look OK. But when I clicked "initialize" the Total Collimation Values changed to yellow & although the Mirror Chase is "Idle", the Platform Position Z = -0.02 and Z(mm) Total Collimation should be 0.773. (for focusoff = 0). The Z=-0.02 was for the -0.8 mm extrafocal, but DX Z is not moving back to the 0mm focus offset. Oops When David clicked "Collimate", the mirror moved back to normal. DX doesn't like 'remote' commands?

23:30 Cleared Active Optics again, running dofpia again.

23:35 Closing & Waiting: David says there is one thick cloud above us, so thick it is blocking the moon even, so we are closing for the moment. It is clear around this cloud so we may be able to open again later. What is this type of cloud called? Fantosi's cloud?

00:15 Re-opening. Talked to Andrew about difficult start with Red. Because of L2 adjustment we needed to change the focus in the collimation lookup table. I edited the DXPMLBCCollimation.dat table to have a Z global offset = -1.6mm (M1 needs to move away from hub by 1.6mm).

00:20 Starting dofpia again. Red pupils are bigger now - more like they should be. But they have coma and FPIA didn't do very well. R 072635 is the image number. On next image 072927, dofpia crashed. Asked David to give Z7 = +2000nm to DX.

Not enough points for estimating Z22
Rejected pupil 0 with diameter 815.449

This manual coma correction helped. Now dofpia is running on these. We probably need to re-do the DX collimation lookup table, given the changes (though I don't know why this affected coma). DX temp is 12.95 C - reported by PSF GUI, but ambient is 9.8 C. Mirror Ventilation is running, but we were shut all last night and all day until opening late.

00:39 Red and Blue converged (time between iterations seemed longer than usual - no hard evidence, just seemed that way). Now running rV superfoc sequence... Starts with 074156(R) and 074158(B). PA=0. Took one at PA=180 again for kicks - while downloading the PA=0 sequences.

00:54 Looks to me like the best focus is between images 1 and 2 of the 5-image sequence at r-SLOAN - so around 0.0mm instead of the nominal best-focus of 0.057. So I guess also the change of L2 will affect the absolute R focus, and therefore all filter focus offsets need to be adjusted accordingly.

01:00 Talked to Andrew - he doesn't think the adjustment of L2 should have affected focus, but is checking. We're concerned that the best-focus offset may be due to instabilities in temperature, dome seeing after having just re-opened. Note that we see clear astigmatism. Will check pointing and test science OB while waiting for things to stabilize a little more.
Note from JMH the following day: I concur that changing L2 should not change the filter focus offset, but residual Z11/Z22 problems from the mirror being out of temperature equilibirum certainly will affect it.

01:20 Going to ACT0276. dofpia here. Wow - FPIA says B seeing is about 2.1" but R seeing is about 0.8". Blue converged pupil image is 083330. Also on R side there is an oscillation in Z7: 083328 to 083529.

After 7 iterations, R was not converging - blue/red largest values on last pupil were ~640nm... so I stopped with "S".

R pointing image: 084227. Old CA, IE were -28.95, -23.17. New CA, IE were -56.5, -74.6 --- a large change, this time in both az and el. We are using old mcspu and old TCS, build 29 (not 31 as on 20090902). And reading az strip encoders again.

01:52 Getting ready to take another superfoc sequence - have yet to run a science OB. dofpia was slow to converge on the pointing star. Converged - now FPIA reports estimated seeing of 0.44"(R) and 1.12"(B). *But, no... R did not converge. Last Z22 = -100nm, which is just the limit. So while B had converged, R had not and so the sequence continued. See /home/LBTO/lbcfpiaRedLog and lbcfpiaBlueLog and images around 085446-090427. After the sets of converged values, for pupils 085825(R) and 085828(B), we see a huge change in focus on R (Z4 = -8745 and next image Z4=8589 and then another set of Z4 oscillations). *

02:00 David giving Z22= +100 to DX. This is needed because about 6 iterations it sent -100nm. We want to remove that.

02:09 Restarting dofpia. Note that with elevation=73deg and T=10.96 (PSF) we have Z=-2.399 for the extra-focal pupil. (dZ = -0.8mm from best-focus). Getting close to limit.

~03:00 Called Andrew - Thinking that we might need to recalculate Z22 zeropoint after L2 was moved, he advised moving L2 to its original position. But he said really it should not change. However, we were going to try to move L2 to the original position to confirm.

I tried to move L2, but I found that the mm position ttyR10 was -6.231 (not -5.98mm as John reported in the afternoon - did it move on it's own or does it not necessarily report the same value?). I tried to move L2 by -400,000 --- each time I found POS = -397 but no motion. Giving up on this.
Note from JMH: The -6.231mm absolute Z2 position was where I had left it. The starting point of these adjustments was not an integer mm.

03:30 Back to dofpia --- again, Red has a radial intensity gradient and Z4 estimation cannot cope. Stopped FPIA and David is clearing active optics.

03:40 Running dofpia again. 104128(B), 104126(R)

03:49 dofpia converged. Running a rV superfoc sequence. Red images are: 104912 - 105130. Now I measure best-focus at center of Chip 2 as 0.047mm (fwhm = 2.65). Looks closer to the nominal best-focus for r-SLOAN (0.057) than the beginning-of-night sequences. Top Chip 2 ~0.04mm and Bottom Chip 2 ~0.05mm

03:59 Running dofpia again. Running superfoc rV again, this time PA=180. R sequence starts with 110045... measure best-focus at center of Chip 2 as 0.038mm (fwhm = 2.34). 0.017mm at top of chip 2

04:21 After another dofpia took a pair of Extra/Intra focal pupil images using all 4 chips. These are 112152 and 112244

04:26 Slewing to a field to test rotator/guiding. 3 images, 120sec each, at U-Bessel and z-SLOAN started at 04:35. This is a random field at 3h, 50deg. Starts with 113532 (R) and 113538 (B). Guiding is working for both blue & red. Blue Guide star is not in focus - Red guide star is. Likewise, B image is not in very good focus (5.5 pixels) but R image is (2.2 pix FWHM).

04:51 Ran dofpia again and now taking a rV superfoc sequence. R images: 115143,115252,...

05:00 Again, after running dofpia taking another rV superfoc sequence. R images: 120048, 120122, 120157, 120232 =We lost an LBC Red image here, the 2nd one of the sequence, i.e. the one between 120048 and 120122.

05:07 dofpia and then extra/intra focal pair. 120726 and 120818

05:14 dofpia and another superfoc sequence: R images are 121449, 121523,..., 121706.

05:20 Closing

Summary

  • Red focal plane curvature is much less after John moved L2 yesterday. (Though much written about trying to move L2 during this night, I could not move it, so it remained where John left it).
  • Rotator Tracking and Guiding seem to be fine with TCS built 29
  • Dan Cox called about uploading the new mcspu, but at that moment we were still not finished with the LBC tests and I was puzzling over why L2 would not move. This test was deferred.
  • I noticed one missing LBC Red image (the one between 120048 and 120122).
  • The DX Collimation lookup table, in /home/telescope/TCS/Configuration/PSF/new/ DXPMLBCCollimation.dat was edited to change Z Global Offset from 0.0 to -1.6mm. This was needed because of the L2 move.

Outstanding Issues

  • why could I not move L2?

  • why is position reported as -6.231mm and not the -5.98 mm at which John left it yesterday?
    JMH: -6.231 is where I left it. -5.98 was how much I moved it.

  • FPIA seems to be running slowly. Could /Repository be mounted in the way that /TestRepo was last June - that really helped.
    JMH: /TestRepo was still mounted, but your LBCFPIA_DATADIR variable was set to /Repository, so I have changed it at 17:50 MST on 4-September.

  • Why does best-focus change - from start of night to middle of night, e.g. Is it all due to being out of thermal equilibrium. Ambient-glass temp difference was high early on, about 4C, but later got as low as 0.5C but towards morning increased as ambient temp dropped.
    JMH: Yes, having the glass 4 degC out of equilibrium is enough to change the best focus because of residual Z11 and Z22. Although, the effect is caused by how fast the mirror is changing temperature rather than how far away from ambient temperature it is.

  • Does Z22 zero point need to be recalibrated (or Z11) after L2 move? Think not...
    JMH: They may need recalibrated based on images Andrew was showing me yesterday, but not because of the L2 move.

-- OlgaKuhn - 04 Sep 2009
Topic revision: r3 - 05 Sep 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback